OpenStudy (anonymous):

passing through (-2, 9/4), vertx at (4, 0).. find the equation of ellipse.. center @ origin..

OpenStudy (anonymous):

what

OpenStudy (anonymous):



OpenStudy (anonymous):

let the eq. of ellipse be : \[\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}=1\],a>b

OpenStudy (anonymous):

yes i substituted the points on the equation and viola.. stuck up

OpenStudy (anonymous):

therefore a=4. Now since (-2,9/4) lies on it ,therefore, \[\frac{(-2)^{2}}{4^{2}}+\frac{(9/4)^{2}}{b^{2}}=1\] find b.

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

its 3 square root 3 / 2

OpenStudy (anonymous):

a= 4?

OpenStudy (anonymous):

yes since distance between center and vertex is a ,so a=4. and you'll get b^2 = 27/4

OpenStudy (anonymous):

yes!! finally thank you

OpenStudy (anonymous):

:)